Stuck on this one any help would be appreciated!
If -4(x-1)=y^2 and 1=(((x-4)^2)/9)+((y^2)/4), how many solutions does this system have?
Answer is 1 solution

Answer by Alan3354(69443)

If -4(x-1)=y^2 and 1=(((x-4)^2)/9)+((y^2)/4), how many solutions does this system have?
-----------
(x-4)^2/9)+ y^2/4 = 1
Sub for y^2
(x-4)^2/9) -4(x-1)/4 = 1
(x-4)^2/9) -x + 1 = 1
(x-4)^2 - 9x = 0
x^2 - 17x + 16 = 0
(x-16)*(x-1) = 0
x = 1
y^2 = -4(x-1)--> y = 0
Solution (1,0)
--------
x = 16
y^2 = -4(x-1)--> y^2 = -60
Not a solution.